# FAB-PPI: Frequentist, Assisted by Bayes, Prediction-Powered Inference

### v1

@@ 6 Discussion and extensions
We proposed FAB-PPI as a Bayes-informed method to significantly improve the performance of PPI in the presence of high quality predictions. In doing so, we showed that the horseshoe represents a sensible default prior for FAB-PPI, contrary to the seemingly natural choice of a Gaussian prior. However, several options may be worth exploring.
In particular, the horseshoe prior was chosen due to its popularity and key properties: (i) its spike at zero (ii) power-law tails and (iii) the closed-form expression for the marginal density $\pi(y)$. However, many other scale-mixture of Gaussians models share these properties. For example, the family of priors with a beta prime (aka inverted beta) prior over the variance (Polson & Scott, 2012), which includes the horseshoe, normal-exponential-gamma (Griffin & Brown, 2011) and other robust priors (Berger, 1980; Strawderman, 1971) as special cases, share the same three properties. On the other hand, some other standard priors such as the Laplace prior (Park & Casella, 2008) or normal-gamma prior (Caron & Doucet, 2008; Griffin & Brown, 2010) do not have power-law tails and therefore do not offer the same robustness guarantees. Other priors, such as the student-t, lack an analytical expression for $\pi(y)$, therefore requiring additional numerical approximation to be applied to FAB-PPI.
@@
As a potential drawback, FAB-PPI shares the computational limitations of the PPI approach (Angelopoulos et al., 2023b), which are discussed in Angelopoulos et al. (2023a). In particular, except for particular cases such as mean estimation and linear regression, the method requires evaluating $\widehat{m}_{\theta}+\widehat{\Delta}_{\theta}$ over a grid of values of $\theta$. This can be computationally expensive, especially in high-dimensional settings.
@@ Acknowledgements
where $\eta>0$ and $C^{+}(0,1)$ is the half-Cauchy distribution with location parameter 00 and scale parameter $1$. Throughout this section and the main text, we assume $\eta=1$. The rationale for this choice, along with a discussion of the general case $\eta\neq 1$, is provided at the end of this section.
@@
[…]
@@
Applying Tweedie’s formula , we obtain the posterior mean
@@
In this section and in the main text, we focused on the specific parametrisation $\eta=1$. For a general $\eta$, similar expressions can be derived for the marginal likelihood and posterior mean, replacing Kummer’s ${}_{1}F_{1}$ function with the more general degenerate hypergeometric function of two variables, $\Phi_{1}$ (see (Carvalho et al., 2010, Equations (4) in the main text and (A1) in the appendix)). While Kummer’s ${}_{1}F_{1}$ function is implemented in many standard scientific libraries, such as SciPy, $\Phi_{1}$ is not. Consequently, computing the marginal likelihood when $\eta\neq 1$ requires numerical integration. Since the evaluation of the marginal likelihood is crucial to our approach, it is therefore reasonable to set $\eta=1$ here.
@@
[…]
@@
[…]
@@
[…]
@@
$H_{0}:\beta=\beta^{\prime}\text{ vs }H_{1}:\beta\sim\pi_{0}$
@@
where $k_{\alpha}(\beta^{\prime})$ is such that $\Pr(Y\in A_{\alpha}(\beta)\mid\beta=\beta^{\prime})=1-\alpha$. The acceptance region is an interval $[\underline{A}_{\alpha}(\beta^{\prime}),\overline{A}_{\alpha}(\beta^{\prime})]$ (Cortinovis & Caron, 2024, Theorem 2.1). Defining
@@
$\mathcal{C}_{\alpha}(y)=\{\beta^{\prime}\mid\underline{A}_{\alpha}(\beta^{\prime})=\beta^{\prime}-\sigma z_{1-\alpha w_{\alpha}(\beta^{\prime})}\leq y\leq\beta^{\prime}+\sigma z_{1-\alpha(1-w_{\alpha}(\beta^{\prime}))}=\overline{A}_{\alpha}(\beta^{\prime})\}.$
The function $w_{\alpha}(\beta^{\prime})\in[0,1]$ is called the spending function or tail function (Puza & O’Neill, 2006; Yu & Hoff, 2018), and represents the proportion of the $\alpha$ rejection budget allocated to the left tail of the acceptance interval $[\underline{A}_{\alpha}(\beta^{\prime}),\overline{A}_{\alpha}(\beta^{\prime})]$.
@@
If the prior $\pi_{0}(\beta)=\mathcal{N}(\beta;0,\tau^{2}\sigma^{2})$ is Gaussian, the spending function is given by $w_{\alpha}(\beta)=g_{\alpha}^{-1}\left(\frac{2\beta}{\sigma\tau^{2}}\right)$ where $g_{\alpha}:(0,1)\to\mathbb{R}$ is the one-to-one function
@@
Proposition S1.3 (FAB with a prior with power-law tails (Cortinovis & Caron, 2024)).
Let $\pi_{0}(\beta;\sigma)$ be a symmetric prior such that the marginal density $\pi(y)$ has power-law tails:
$\pi(y)\sim\frac{C_{\sigma}}{|y|^{2\beta+1}}\text{ as }|y|\to\infty$
for some constant $C_{\sigma}$ and some $\beta>0$. Then,
@@
Here, we outline the steps to derive the FAB-PPI mean estimator presented in Equation 24, as well as the corresponding FAB-PPI confidence region. The convex loss function that corresponds to estimating $\theta^{*}=\mathbb{E}[Y]$ is the squared loss $\mathcal{L}_{\theta}(x,y)=\frac{1}{2}\left(\theta-y\right)^{2}$. In this case, the subgradient of $\mathcal{L}_{\theta}$ with respect to $\theta$ is given by $\mathcal{L}_{\theta}^{\prime}(x,y)=\theta-y$. As a result of this, the measure of fit $m_{\theta}$ and the rectifier $\Delta_{\theta}$ take the form
@@
$\Delta_{\theta}$ $=\mathbb{E}[\mathcal{L}_{\theta}(X,Y)-\mathcal{L}_{\theta}^{\prime}(X,f(X))]=\mathbb{E}[f(X)-Y].$
@@
$\widehat{m}_{\theta}=\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}_{\theta}^{\prime}(X_{i},f(X_{i}))=\theta-\frac{1}{N}\sum_{i=1}^{N}f(\widetilde{X}_{i}),$
@@
The sample mean estimator (10) and the control variate estimator (12) under the squared loss are recovered by setting $\lambda$ to $1$ and $\widehat{\lambda}$ (14), respectively. From this, the standard PPI mean estimators (11) and (13) are obtained by solving the equation $\widehat{m}_{\theta}+\widehat{\Delta}=0$ for $\theta$.
Instead, we first define the Bayes assisted estimator (20) under the chosen prior $\pi_{0}(\Delta;\tau_{n})$,
@@
where $\widehat{\sigma}^{2}$ is an estimator of $\mathrm{var}(\widehat{\Delta})$ and $\ell_{\theta}^{\prime}(z;\sigma,\tau)$ is the derivative of the log-marginal likelihood of a Gaussian likelihood model with mean $\Delta$ and variance $\sigma^{2}$ under the prior $\pi_{0}(\Delta,\tau)$. Then, the FAB-PPI mean estimator $\widehat{\theta}^{\mathrm{\texttt{FABPP}}}$ under $\pi_{0}$ is given by the solution to the equation
@@
$\mathcal{T}_{\alpha-\delta}(\widehat{m}_{\theta})=(\widehat{m}_{\theta}\pm\widehat{\sigma}^{f}z_{1-(\alpha-\delta)/2})=\left(\theta-\widehat{\theta}^{f}\pm\widehat{\sigma}^{f}z_{1-(\alpha-\delta)/2}\right),$
@@
where, again, $\widehat{\sigma}^{2}$ is an estimator of $\mathrm{var}(\widehat{\Delta})$. Finally, to avoid making assumptions on the specific form of $\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta}(\widehat{\Delta})$, we use $(\inf\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta}),\sup(\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta}))\supseteq\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta}(\widehat{\Delta})$ in the definition of $\mathcal{C}_{\alpha}^{\mathrm{\texttt{FABPP}}}$ (21) to obtain the FAB-PPI interval
$\mathcal{C}_{\alpha}^{\mathrm{\texttt{FABPP}}}$ =$=\left\{\theta\mid 0\in\left(\theta-\widehat{\theta}^{f}\pm\widehat{\sigma}^{f}z_{1-(\alpha-\delta)/2}\right)+(\inf(\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta}),\sup(\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta}))\right\}$
=$=\left\{\theta\mid 0\in\left(\theta-\widehat{\theta}^{f}-\widehat{\sigma}^{f}z_{1-(\alpha-\delta)/2}+\inf(\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta}),\theta-\widehat{\theta}^{f}+\widehat{\sigma}^{f}z_{1-(\alpha-\delta)/2}+\sup(\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta})\right)\right\}$
$=\left(\widehat{\theta}^{f}-\widehat{\sigma}^{f}z_{1-(\alpha-\delta)/2}-\sup(\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta}),\widehat{\theta}^{f}+\widehat{\sigma}^{f}z_{1-(\alpha-\delta)/2}-\inf(\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta})\right).$
@@
  $(\widehat{\sigma}^{f})^{2}\leftarrow\frac{1}{N-1}\sum_{j=1}^{N}(f(\widetilde{X}_{j})-\widehat{\theta}^{f})^{2}$
@@
  $\widehat{\sigma}^{2}\leftarrow\frac{1}{n}\widehat{\sigma}_{\xi}^{2}+\frac{(\widehat{\lambda}-1)^{2}}{N}(\widehat{\sigma}^{f})^{2}$
@@
Outputs: estimator $\widehat{\theta}^{\mathrm{\texttt{FABPP}}}=\widehat{\theta}^{f}-\widehat{\Delta}-\widehat{\sigma}^{2}\ell^{\prime}(\widehat{\Delta};\widehat{\sigma},\tau_{n})$ and CR $\mathcal{C}_{\alpha}^{\mathrm{\texttt{FABPP}}}=(\widehat{\theta}^{f}-\sigma^{f}z_{1-(\alpha-\delta)/2}-\sup(\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta}),\widehat{\theta}^{f}+\sigma^{f}z_{1-(\alpha-\delta)/2}-\inf(\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta}))$
@@
$\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta}(\widehat{\Delta}_{\theta};\widehat{\sigma}_{\theta})=\{\Delta_{\theta}\mid\Delta_{\theta}-\widehat{\sigma}_{\theta}z_{1-\delta w_{\delta}(\Delta_{\theta};\widehat{\sigma}_{\theta})}<\widehat{\Delta}_{\theta}<\Delta_{\theta}+\widehat{\sigma}_{\theta}z_{1-\delta(1-w_{\delta}(\Delta_{\theta};\widehat{\sigma}_{\theta}))}\},$ (S32)
where $w_{\delta}(\cdot;\widehat{\sigma}_{\theta})$ is the FAB spending function. The proof of asymptotic coverage is organised as follows. First, we show that $\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta}(\widehat{\Delta}_{\theta};\widehat{\sigma}_{\theta})$ is asymptotically a $1-\delta$ confidence interval. This is established via Lemma S3.1 for the Gaussian prior, and via Lemmas S3.2 and S3.3 for the horseshoe prior. This result is then combined with the asymptotic coverage of the standard sample mean estimator for $m_{\theta}$ to conclude the asymptotic coverage of the FAB-PPI estimator of $\theta^{\star}$.
@@
${\lim}_{\min(n,N)\to\infty}\Pr(\Delta_{\theta}\in\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta}(\widehat{\Delta}_{\theta};\widehat{\sigma}_{\theta})\mid\Delta_{\theta})\geq 1-\delta.$ (S33)
@@
Using Proposition S1.2, for any $x>0$, $\sigma>0$,
@@
$\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta}(\widehat{\Delta}_{\theta};\widehat{\sigma}_{\theta})$ =$=\{\Delta_{\theta}>0\mid\Delta_{\theta}-\widehat{\sigma}_{\theta}z_{1-\delta w_{\delta}(\Delta_{\theta};\widehat{\sigma}_{\theta})}<\widehat{\Delta}_{\theta}<3\Delta_{\theta}+\widehat{\sigma}_{\theta}z_{1-\delta w_{\delta}(\Delta_{\theta};\widehat{\sigma}_{\theta})}\}$
$~{}~{}~{}\cup\{\Delta_{\theta}<0\mid 3\Delta_{\theta}-\widehat{\sigma}_{\theta}z_{1-\delta(1-w_{\delta}(\Delta_{\theta};\widehat{\sigma}_{\theta}))}<\widehat{\Delta}_{\theta}<\Delta_{\theta}+\widehat{\sigma}_{\theta}z_{1-\delta(1-w_{\delta}(\Delta_{\theta};\widehat{\sigma}_{\theta}))}\}.$
Additionally, for any $x>0$,
@@
It follows from Equation S32 that, for any $z_{1-\delta}>\epsilon>0$, there exist $N_{0}$ such that for all $n,N$ with $\min(n,N)\geq N_{0}$, the FAB confidence region $\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta}(\widehat{\Delta}_{\theta};\widehat{\sigma}_{\theta})$ contains the set
$\mathcal{S}_{\delta}(\widehat{\Delta}_{\theta};\widehat{\sigma}_{\theta})$ =$=\left\{\Delta_{\theta}>0\mid\Delta_{\theta}-\widehat{\sigma}_{\theta}(z_{1-\delta}-\epsilon)<\widehat{\Delta}_{\theta}<3\Delta_{\theta}+\widehat{\sigma}_{\theta}(z_{1-\delta}-\epsilon)\right\}$
$~{}~{}~{}\cup\left\{\Delta_{\theta}<0\mid 3\Delta_{\theta}-\widehat{\sigma}_{\theta}(z_{1-\delta}-\epsilon)<\widehat{\Delta}_{\theta}<\Delta_{\theta}-\widehat{\sigma}_{\theta}(z_{1-\delta}-\epsilon)\right\}.$

### v2

@@ 6 Discussion and Extensions
We proposed FAB-PPI as a Bayes-informed method to significantly improve the performance of PPI in the presence of high-quality predictions. In doing so, we showed that the horseshoe represents a sensible default prior for FAB-PPI, contrary to the seemingly natural choice of a Gaussian prior. However, several options may be worth exploring.
In particular, the horseshoe prior was chosen due to its popularity and key properties: (i) its spike at zero (ii) power-law tails and (iii) the closed-form expression for the marginal density $\pi(y)$. However, many other scale-mixture of Gaussians models share these properties. For example, the family of priors with a beta prime (aka inverted beta) prior over the variance (Polson & Scott, 2012), which includes the horseshoe, normal-exponential-gamma (Griffin & Brown, 2011) and other robust priors (Berger, 1980; Strawderman, 1971) as special cases, shares the same three properties. On the other hand, some other standard priors such as the Laplace prior (Park & Casella, 2008) or normal-gamma prior (Caron & Doucet, 2008; Griffin & Brown, 2010) do not have power-law tails and therefore do not offer the same robustness guarantees. Other priors, such as the Student-t, lack an analytical expression for $\pi(y)$, therefore requiring additional numerical approximation to be applied to FAB-PPI.
@@
As a potential drawback, FAB-PPI shares the computational limitations of the PPI approach (Angelopoulos et al., 2023a), which are discussed in Angelopoulos et al. (2023b). In particular, except for special cases such as mean estimation and linear regression, the method requires evaluating $\widehat{m}_{\theta}+\widehat{\Delta}_{\theta}$ over a grid of values of $\theta$. This can be computationally expensive, especially in high-dimensional settings.
Supplementary Material and Code.
The supplementary material contains additional background, proofs, and experiments. All sections, figures, and equations in the supplementary material are prefixed with ‘S’ for clarity. Code for reproducing the experiments is available at https://github.com/stefanocortinovis/fab-ppi.
@@ Acknowledgements
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@
where $\eta>0$ and $C^{+}(0,1)$ is the half-Cauchy distribution with location parameter 0 and scale parameter $1$. Throughout this section and the main text, we assume $\eta=1$. The rationale for this choice, along with a discussion of the general case $\eta\neq 1$, is provided at the end of this section.
@@
$\pi(y)$ $=\int_{-\infty}^{\infty}\mathcal{N}(y\mid\beta,\sigma^{2})\pi_{\mathrm{HS}}(\beta)d\beta$
$=\frac{1}{\sqrt{2\pi\sigma^{2}}}\int_{0}^{\infty}\frac{1}{\sqrt{1+\nu^{2}}}e^{-\frac{y^{2}}{2\sigma^{2}(1+\nu^{2})}}p(\nu)d\nu$
$=\frac{2}{\pi\sqrt{2\pi\sigma^{2}}}\int_{0}^{\infty}e^{-\frac{y^{2}}{2\sigma^{2}(1+\nu^{2})}}\frac{1}{(1+\nu^{2})^{3/2}}d\nu.$
@@
Applying Tweedie’s formula (Efron, 2011), we obtain the posterior mean
@@
In this section and in the main text, we focused on the specific parameterisation $\eta=1$. For a general $\eta$, similar expressions can be derived for the marginal likelihood and posterior mean, replacing Kummer’s $_{1}F_{1}$ function with the more general degenerate hypergeometric function of two variables, $\Phi_{1}$ (see (Carvalho et al., 2010, Equations (4) in the main text and (A1) in the appendix)). While Kummer’s $_{1}F_{1}$ function is implemented in many standard scientific libraries, such as SciPy, $\Phi_{1}$ is not. Consequently, computing the marginal likelihood when $\eta\neq 1$ requires numerical integration. Since the evaluation of the marginal likelihood is crucial to our approach, it is therefore reasonable to set $\eta=1$ here.
@@
In this section, we provide additional background on the FAB framework (Pratt, 1961, 1963; Yu & Hoff, 2018).
Let $Y\mid\beta\sim\mathcal{N}(\beta,\sigma^{2})$ with some prior $\pi_{0}(\beta)$. Denote by $\pi(y)=\int_{\mathbb{R}}p(y\mid\beta)\pi_{0}(\beta)d\beta$ the corresponding marginal likelihood. For $\alpha\in(0,1)$, let $\mathcal{C}_{\alpha}$ be the confidence procedure that solves the constrained optimisation problem
@@
where $\mathrm{vol}(\mathcal{C}_{\alpha}(y))=\int_{\beta^{\prime}\in\mathcal{C}_{\alpha}(y)}d\beta^{\prime}$ is the volume of $\mathcal{C}_{\alpha}(y)$ and
$\mathbb{E}[\mathrm{vol}(\mathcal{C}_{\alpha}(Y))]=\int_{\mathbb{R}}\mathrm{vol}(\mathcal{C}_{\alpha}(y))\pi(y)dy$ (S28)
@@
$\mathbb{E}[\mathrm{vol}(\mathcal{C}_{\alpha}(Y))]$ $=\int_{\mathbb{R}}\mathrm{vol}(\mathcal{C}_{\alpha}(y))\pi(y)dy$
$=\int_{\mathbb{R}}\int_{\mathbb{R}}1_{\beta^{\prime}\in\mathcal{C}_{\alpha}(y)}d\beta^{\prime}\pi(y)dy$
$=\int_{\mathbb{R}}\Pr(\beta^{\prime}\in\mathcal{C}_{\alpha}(Y))d\beta^{\prime}.$
@@
$H_{0}:\beta=\beta^{\prime}\text{ vs }H_{1}:\beta\sim\pi_{0},$
@@
where $k_{\alpha}(\beta^{\prime})$ is such that $\Pr(Y\in A_{\alpha}(\beta)\mid\beta=\beta^{\prime})=1-\alpha$. The acceptance region is an interval $\underline{A}_{\alpha}(\beta^{\prime}),\overline{A}_{\alpha}(\beta^{\prime})]$ (Cortinovis & Caron, 2024, Theorem 3.3). Defining
@@
$\mathcal{C}_{\alpha}(y)=\{\beta^{\prime}\mid\underline{A}_{\alpha}(\beta^{\prime})=\beta^{\prime}-\sigma z_{1-\alpha w_{\alpha}(\beta^{\prime})}\leq y\leq\beta^{\prime}+\sigma z_{1-\alpha(1-w_{\alpha}(\beta^{\prime}))}=\overline{A}_{\alpha}(\beta^{\prime})\}.$ (S29)
The function $w_{\alpha}(\beta^{\prime})\in[0,1]$ is called the spending function or tail function (Puza & O’Neill, 2006; Yu & Hoff, 2018), and represents the proportion of the $\alpha$ rejection budget allocated to the left tail of the acceptance interval $\underline{A}_{\alpha}(\beta^{\prime}),\overline{A}_{\alpha}(\beta^{\prime})]$.
@@
If the prior $\pi_{0}(\beta)=\mathcal{N}(\beta;0,\tau^{2}\sigma^{2})$ is Gaussian, the spending function is given by $w_{\alpha}(\beta)=g_{\alpha}^{-1}\left(\frac{2\beta}{\sigma\tau^{2}}\right)$, where $g_{\alpha}:(0,1)\to\mathbb{R}$ is the one-to-one function
@@
Proposition S1.3 (FAB with a prior with power-law tails (Cortinovis & Caron, 2024, Lemma S1.1)).
Let $\pi_{0}(\beta;\sigma)$ be a symmetric prior on $\beta$ such that the marginal density $\pi(y)$ has power-law tails, i.e.
$\pi(y)\sim C_{\sigma}|y|^{-\delta}\text{ as }|y|\to\infty$
for some constant $C_{\sigma}$ and some exponent $\delta>1$. Then, $w_{\alpha}(\beta;\sigma)$ is bounded away from 0 and $1$, and
@@
Here, we outline the steps to derive the FAB-PPI mean estimator presented in Equation 24, as well as the corresponding FAB-PPI confidence region.
The convex loss function that corresponds to estimating $\theta^{\star}=\mathbb{E}[Y]$ is the squared loss $\mathcal{L}_{\theta}(x,y)=\frac{1}{2}\left(\theta-y\right)^{2}$. In this case, the subgradient of $\mathcal{L}_{\theta}$ with respect to $\theta$ is given by $\mathcal{L}_{\theta}^{\prime}(x,y)=\theta-y$. As a result of this, the measure of fit $m_{\theta}$ and the rectifier $\Delta_{\theta}$ take the form
@@
$\Delta_{\theta}$ $=\mathbb{E}[\mathcal{L}_{\theta}^{\prime}(X,Y)-\mathcal{L}_{\theta}^{\prime}(X,f(X))]=\mathbb{E}[f(X)-Y].$
@@
$\widehat{m}_{\theta}=\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}_{\theta}^{\prime}(\widetilde{X}_{i},f(\widetilde{X}_{i}))=\theta-\frac{1}{N}\sum_{i=1}^{N}f(\widetilde{X}_{i}),$
@@
The sample mean estimator (10) and the control variate estimator (12) under the squared loss are recovered by setting $\lambda$ to $1$ and $\widehat{\lambda}$ as in Equation 14, respectively. From this, the standard PPI mean estimators (11) and (13) are obtained by solving the equation $\widehat{m}_{\theta}+\widehat{\Delta}=0$ for $\theta$.
Instead, we first define the Bayes-assisted estimator (20) under the chosen prior $\pi_{0}(\Delta;\tau_{n})$,
@@
[…]
@@
$\mathcal{T}_{\alpha-\delta}(\widehat{m}_{\theta})=[\widehat{m}_{\theta}\pm\widehat{\sigma}^{f}z_{1-(\alpha-\delta)/2}]=\left[\theta-\widehat{\theta}^{f}\pm\widehat{\sigma}^{f}z_{1-(\alpha-\delta)/2}\right],$
@@
[…]
$\mathcal{C}_{\alpha}^{\mathrm{\texttt{FABPP}}}$ $=\left\{\theta\mid 0\in\left[\theta-\widehat{\theta}^{f}\pm\widehat{\sigma}^{f}z_{1-(\alpha-\delta)/2}\right]+[\inf(\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta}),\sup(\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta})]\right\}$
$=\left\{\theta\mid 0\in\left[\theta-\widehat{\theta}^{f}-\widehat{\sigma}^{f}z_{1-(\alpha-\delta)/2}+\inf(\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta}),\theta-\widehat{\theta}^{f}+\widehat{\sigma}^{f}z_{1-(\alpha-\delta)/2}+\sup(\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta})\right]\right\}$
$=\left[\widehat{\theta}^{f}-\widehat{\sigma}^{f}z_{1-(\alpha-\delta)/2}-\sup(\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta}),\widehat{\theta}^{f}+\widehat{\sigma}^{f}z_{1-(\alpha-\delta)/2}-\inf(\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta})\right].$
@@
$(\widehat{\sigma}^{f})^{2}\leftarrow\frac{1}{N(N-1)}\sum_{j=1}^{N}(f(\widetilde{X}_{j})-\widehat{\theta}^{f})^{2}$
@@
$\widehat{\sigma}^{2}\leftarrow\frac{1}{n}\widehat{\sigma}_{\xi}^{2}+(\widehat{\lambda}-1)^{2}(\widehat{\sigma}^{f})^{2}$
@@
Outputs: estimator $\widehat{\theta}^{\mathrm{\texttt{FABPP}}}=\widehat{\theta}^{f}-\widehat{\Delta}-\widehat{\sigma}^{2}\ell^{\prime}(\widehat{\Delta};\widehat{\sigma},\tau_{n})$ and CR $\mathcal{C}_{\alpha}^{\mathrm{\texttt{FABPP}}}=[\widehat{\theta}^{f}-\widehat{\sigma}^{f}z_{1-(\alpha-\delta)/2}-\sup(\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta}),\widehat{\theta}^{f}+\widehat{\sigma}^{f}z_{1-(\alpha-\delta)/2}-\inf(\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta})]$
@@
Some of the results discussed in this section, such as Section S3.1 and Section S3.2, are concerned with the convergence of closed sets with respect to the Hausdorff distance, which we recall here for completeness. Given two closed subsets $C_{1}$ and $C_{2}$ of $\mathbb{R}$, their Hausdorff distance $d_{\mathrm{H}}$ is defined as
$d_{\mathrm{H}}(C_{1},C_{2})=\max\left\{\sup_{x\in C_{1}}\inf_{y\in C_{2}}|x-y|,\sup_{y\in C_{2}}\inf_{x\in C_{1}}|x-y|\right\}.$
Then, for a collection of closed subsets $(C_{1}(y))_{y\in\mathbb{R}}$ and a closed subset $C_{2}$ of $\mathbb{R}$, $(C_{1}(y))_{y\in\mathbb{R}}$ converges in Hausdorff distance to $C_{2}$ if $\lim_{y\to\infty}d_{\mathrm{H}}(C_{1}(y),C_{2})=0$. In particular, if $C_{2}=[a,b]$ is a closed interval for some $a<b$, $\lim_{y\to\infty}d_{\mathrm{H}}(C_{1}(y),C_{2})=0$ if and only if, for all $\epsilon\in(0,\frac{b-a}{2})$, there exists $y_{0}$ such that $a+\epsilon,b-\epsilon]\subseteq C_{1}(y)\subseteq[a-\epsilon,b+\epsilon]$ for all $y>y_{0}$. In the sequel, we write $\lim_{y\to\infty}C_{1}(y)=C_{2}$ for $\lim_{y\to\infty}d_{\mathrm{H}}(C_{1}(y),C_{2})=0$.
@@
$\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta}(\widehat{\Delta}_{\theta};\widehat{\sigma}_{\theta})=\{\Delta_{\theta}\mid\Delta_{\theta}-\widehat{\sigma}_{\theta}z_{1-\delta w_{\delta}(\Delta_{\theta};\widehat{\sigma}_{\theta})}\leq\widehat{\Delta}_{\theta}\leq\Delta_{\theta}+\widehat{\sigma}_{\theta}z_{1-\delta(1-w_{\delta}(\Delta_{\theta};\widehat{\sigma}_{\theta}))}\},$ (S33)
where $w_{\delta}(\cdot;\widehat{\sigma}_{\theta})$ is the FAB spending function. The proof of asymptotic coverage is organised as follows. First, we show that $\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta}(\widehat{\Delta}_{\theta};\widehat{\sigma}_{\theta})$ is asymptotically a $1-\delta$ confidence interval. This is established via Section S3.1 for the Gaussian prior, and via Section S3.1 for the horseshoe prior. This result is then combined with the asymptotic coverage of the standard sample mean estimator for $m_{\theta}$ to conclude the asymptotic coverage of the FAB-PPI estimator of $\theta^{\star}$.
@@
${\lim\inf}_{\min(n,N)\to\infty}\Pr(\Delta_{\theta}\in\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta}(\widehat{\Delta}_{\theta};\widehat{\sigma}_{\theta})\mid\Delta_{\theta})\geq 1-\delta.$ (S34)
@@
Using Section S1.2, for any $x>0$, $\sigma>0$,
@@
$\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta}(\widehat{\Delta}_{\theta};\widehat{\sigma}_{\theta})$ $=\{\Delta_{\theta}>0\mid\Delta_{\theta}-\widehat{\sigma}_{\theta}z_{1-\delta w_{\delta}(\Delta_{\theta};\widehat{\sigma}_{\theta})}\leq\widehat{\Delta}_{\theta}\leq 3\Delta_{\theta}+\widehat{\sigma}_{\theta}z_{1-\delta w_{\delta}(\Delta_{\theta};\widehat{\sigma}_{\theta})}\}$
$~~~\cup\{\Delta_{\theta}<0\mid 3\Delta_{\theta}-\widehat{\sigma}_{\theta}z_{1-\delta(1-w_{\delta}(\Delta_{\theta};\widehat{\sigma}_{\theta}))}\leq\widehat{\Delta}_{\theta}\leq\Delta_{\theta}+\widehat{\sigma}_{\theta}z_{1-\delta(1-w_{\delta}(\Delta_{\theta};\widehat{\sigma}_{\theta}))}\}$
$~~~\cup\{0\mid|\widehat{\Delta}_{\theta}|\leq\widehat{\sigma}_{\theta}z_{1-\delta/2}\}.$
Consider first $\Delta_{\theta}=0$. By the CLT, $\Pr(0\in\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta}(\widehat{\Delta}_{\theta};\widehat{\sigma}_{\theta})\mid\Delta_{\theta}=0)=\Pr(|\widehat{\Delta}_{\theta}|\leq\widehat{\sigma}_{\theta}z_{1-\delta/2}\mid\Delta_{\theta}=0)\to 1-\delta$ as $\min(n,N)\to\infty$. Additionally, for any $x>0$,
@@
It follows from Equation S33 that, for any $\epsilon\in(0,z_{1-\delta})$, there exist $N_{0}$ such that for all $n,N$ with $\min(n,N)\geq N_{0}$, the FAB confidence region $\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta}(\widehat{\Delta}_{\theta};\widehat{\sigma}_{\theta})$ contains the set
$\mathcal{S}_{\delta}(\widehat{\Delta}_{\theta};\widehat{\sigma}_{\theta})$ $=\left\{\Delta_{\theta}>0\mid\Delta_{\theta}-\widehat{\sigma}_{\theta}(z_{1-\delta}-\epsilon)\leq\widehat{\Delta}_{\theta}\leq 3\Delta_{\theta}+\widehat{\sigma}_{\theta}(z_{1-\delta}-\epsilon)\right\}$
$~~~\cup\left\{\Delta_{\theta}<0\mid 3\Delta_{\theta}-\widehat{\sigma}_{\theta}(z_{1-\delta}-\epsilon)\leq\widehat{\Delta}_{\theta}\leq\Delta_{\theta}-\widehat{\sigma}_{\theta}(z_{1-\delta}-\epsilon)\right\}\cup\{0\mid|\widehat{\Delta}_{\theta}|\leq\widehat{\sigma}_{\theta}z_{1-\delta/2}\}.$